OpenStudy (anonymous):

prove that lim x->0 cos(1/x) does not exist.

OpenStudy (agreene):

\[\lim_{x\rightarrow 0}\cos(\frac{1}{x})\] we can start by looking at what 1/x is. If that is undefined then we can assume cos(undefined) = undefined. (or DNE) \[lim_{x \rightarrow 0^+}\frac{1}{x}=\infty \] and \[lim_{x\rightarrow0^-}\frac{1}{x}=-\infty\] therefore \[\lim_{x\rightarrow0}\frac{1}{x}=undefined\] and \[\lim_{x\rightarrow0}\cos(undefined)=undefined\]

OpenStudy (anonymous):

no right and left hand limits yet. can't use them.

OpenStudy (anonymous):

how about this: letx_n = 1/2pi(n), lim x_n = 0 but limf(x_n) = lim cos(2pin) = 1 since lim x_n does not equal lim f (x_n) thr limit does not exist

OpenStudy (anonymous):

the limit does not exist

OpenStudy (agreene):

I think there is a way to show \[\cos\frac{1}{0} = \sin\infty\] and lim x->infty sin(x) = undefined.

OpenStudy (anonymous):

no l'hopitals rule yet either,

OpenStudy (agreene):

Well, that is just annoying.